OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. 3x + 11y = 34.5 9x – 7y = -16.5

OpenStudy (anonymous):

Multiply all the top line by -3 and then add the top line with bottom line. Let me know if you want me to show you how (:

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i did i got x=2 but when i plug it in i get the wrong answer

OpenStudy (anonymous):

just show me please don't know were i went wrong

OpenStudy (anonymous):



OpenStudy (anonymous):

Make sure my calculations are right, Haha (:

OpenStudy (anonymous):

ok thanks i was doing it all wrong so is there mutlp ways of doing this

OpenStudy (anonymous):

so y would be 1.5

OpenStudy (anonymous):

It doesn't matter which one you cancel out. In this case, canceling the x is easier since they have a like multiple (: Like for 11 & 7 you'd have to multiply the top by 7 and the bottom by 11. Wayyy long.

OpenStudy (anonymous):

ok thats what i was trying to do so y=1.5

OpenStudy (anonymous):

x=1/2 & y=3

OpenStudy (anonymous):

So x=.5 not 1.5 and y=3

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

can you show me how to use x

OpenStudy (anonymous):

You just plug the y (3) into the 3x + 11y = 34.5 as 3x + 11(3) = 34.5 and just solve it as a simple equation trying to find x (:

OpenStudy (anonymous):

ok i figured it out thanks